briyamilet1:

Determine the axis of symmetry for the function f(x) = −2(x + 3)2 − 5. x = 3 x = −3 x = 5 x = −5

payload:

So this is a quadratic in the vertex form f(x) = a(x-h)^2+k where the vertex is (h,k) and the axis of symmetry is x=h so here h=-3, right? so then the axis of symmetry is x=h, plug it in

briyamilet1:

so... x=3?

payload:

-3 not 3

briyamilet1:

oh. oops haha. ty :D

payload:

f(x) = a(x-h)^2+k 3=-h so h=-3 (:

briyamilet1:

Thank youuuu